HOW TO SOLVE THIS ON A FINANCIAL CALCULATOR? Burnett Corp. pays a constant $8.25 dividend on its stock. The company will maintain this dividend for the next 13 years and will then cease paying dividends forever. If the required return on this stock is 11.2 percent, what is the current share price? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1
As per dividend discount model, current share price is the present value of future dividends.
Current share price = Present value of future dividends
= Annual dividend * Present value of annuity of 1
= $       8.25 * 6.68252
= $    55.13
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.1120)^-13)/0.1120 i = 11.20%
= 6.68252 n = 13